OpenStudy (anonymous):

If the graphs of 2y +x +3 =0 and 3y +ax +2 =0 are perpendicular, find the value of “a”.

OpenStudy (anonymous):

If the lines are perpendicular their product of slopes = -1 ie m1*m2 = -1

OpenStudy (anonymous):

1 st find the slope of two equation

OpenStudy (anonymous):

@Yahoo! then just multiply the slope of one line by the opposite reciprocal and solve for a, correct?

OpenStudy (anonymous):

2y=-x-3 y=(-1/2)x-3 using equation y=mx+c m1=-1/2 3y=-ax-2 y=(-a/3)x-2 using again y=mx+c m2=-a/3 for perpendicular lines m1xm2=-1 (-1/2)x(-a/3)= -1 a/6=-1 a=-6 is your answer

OpenStudy (anonymous):

Whenever you get there kind of questions remember that product of slopes of two perpendicular lines in always -1